OpenStudy (anonymous):

what is the phase shift of h(t)=-2cos(5t-pi/7)

OpenStudy (anonymous):

A wave is described by an equation in the form of: \[y(t) = Acos(\omega t - \phi)\] where \[A = Amplitude, \omega = Angular Frequency, \phi = Phase Shift\]

OpenStudy (anonymous):

You can also switch in a sin for the cos. The only thing that'll change is the phase shift. But you should be able to pick out the phase shift easily.

OpenStudy (anonymous):

it does not do throw in my my homework system

OpenStudy (anonymous):

i entered pi/7

OpenStudy (anonymous):

do i have change the sign or something???

OpenStudy (anonymous):

Oh, I completely blanked on something. We need to get the function into a slightly different form: \[y = A \cos (\omega t - \phi) \rightarrow A \cos(\omega(t - \frac{\phi}{\omega}))\] Your phase shift would then be phi/omega

OpenStudy (anonymous):

thank you i got it